# Topological Optimal Transport for Geometric Cycle Matching

### v1

@@ 4.3 Geodesics in the space of measure topological networks
$\qquad\qquad=\|\omega_{t}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{t}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{t},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}$
$\hskip 144.54pt+t(1-t)\left(\|\omega_{0}-\omega_{1}\|^{2}_{L^{2}(\pi^{v}\otimes\pi^{e})}+\|k_{0}-k_{1}\|^{2}_{L^{2}(\pi^{v}\otimes\pi^{v})}+\|d_{2}\circ(\iota_{0},\iota_{1})\|^{2}_{L^{2}(\pi^{e})}\right)$
$\qquad\qquad=\|\omega_{t}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{t}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{t},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}$
$\hskip 144.54pt+t(1-t)\left(\|\omega_{0}-\omega_{1}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{0}-k_{1}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{0},\iota_{1})\|^{2}_{L^{2}(\xi^{e})}\right)$ (38)
$\qquad\qquad=(1-t)\left(\|\omega_{0}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{0}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{0},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}\right)$
$\hskip 144.54pt+t\left(\|\omega_{1}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{1}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{1},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}\right)$ (39)
$\qquad\qquad\geq(1-t)d_{\rm TpOT,2}(P^{0},P^{\prime})^{2}+td_{\rm TpOT,2}(P^{1},P^{\prime})^{2},$ (40)
where 38 follows by marginal properties of $\xi^{v}$ and $\xi^{e}$, 39 by definition of $w_{t},k_{t}$ and $\nu_{t}$, and 40 by suboptimality. ∎
5 Examples
In this section, we demonstrate our computational framework and the capabilities of TpOT on a range of numerical examples. In what follows, persistent homology computations are performed using the Julia software Ripserer.jl [14]. Specifically, we compute the Vietoris-Rips filtration and we use the involutive algorithm [14, 15] to compute homology and representatives. We emphasise that the choice of filtration and representatives influences the resulting transport plan and analysis.
Figure 3: (A) Source point cloud: four circles in $\mathbb{R}^{2}$. (B) Target point cloud, with points coloured according to the Gromov-Wasserstein coupling. (C) Target point cloud, coloured according to the TpOT coupling with $\alpha=0.5$ and $\beta=1$. (D) As $\alpha$ varies, TpOT interpolates between a coupling driven by geometry (Gromov-Wasserstein, $\alpha\uparrow 1$) and topological features (persistence diagram matchings, $\alpha\downarrow 0$). (E) Geodesic between the point clouds for $\beta=1$.
@@ 5.1 Matching point clouds
In Figure 3, we consider a source point cloud $X$ consisting of four circles of uniform size (Figure 3(A)) and a target point cloud $X^{\prime}$ resembling a “flower” with four ellipses as “petals” (Figure 3(B)). The point clouds $X$ and $X^{\prime}$, together with the uniform distribution on points, are then measure spaces. We choose to endow these measure spaces with Gaussian kernels $k,k^{\prime}$ so that $k_{ij}=K(x_{i},x_{j})$ and $k^{\prime}_{ij}=K(x_{i}^{\prime},x_{j}^{\prime})$ where
$K(x,y)=\exp\left(-\frac{\|x-y\|_{2}^{2}}{h}\right),$
where the bandwidth $h$ is chosen for each set of points $\{x_{1},\ldots,x_{N}\}$ such that $h^{-1}\left(N^{-2}\sum_{ij}\|x_{i}-x_{j}\|_{2}^{2}\right)=1$.
In Figure 3(B), points of the target point cloud are coloured by their corresponding source points under the Gromov-Wasserstein matching. Clearly, topological features are not preserved by this matching. Instead, each of the circles in the source point cloud is split up among multiple ellipses in the target point cloud.
@@
The $1$-dimensional persistence diagram of the Vietoris-Rips filtration $\mathcal{K}_{\rm VR}(X)$ contains 4 significant, almost identical, homology classes (one per circle), and the same is true for $\mathcal{K}_{\rm VR}(X^{\prime})$ (one class per ellipse). A choice of generating cycle for each class in the diagrams yields the PH-hypergraphs $H_{X}=H(X,\mathcal{K}_{\rm VR}(X),g_{X})$ and $H_{X^{\prime}}=H(X^{\prime},\mathcal{K}_{\rm VR}(X^{\prime}),g_{X^{\prime}})$ and the corresponding topological networks (see Section 3.3)
$P_{X}$ $=\left((X,d_{\mathbb{R}^{n}},\mu_{X}),(Y,\iota,\nu),\omega_{H}\right)$
$P_{X^{\prime}}$ $=\left((X^{\prime},d_{\mathbb{R}^{n}},\mu_{X^{\prime}}),(Y^{\prime},\iota^{\prime},\nu^{\prime}),\omega_{H^{\prime}}\right).$
Using the algorithms we developed in Section 3.3, we are able to simultaneously find matchings between points and topological features in $X$ and $X^{\prime}$ as the parameters $(\alpha,\beta)$ vary.
Figure 4: (A) Source point cloud and target point cloud have almost identical persistence diagrams. (B) Target point cloud, matching induced by GW distance. The colour code represents the matching. (C) Target point cloud, matching induced by TpOT with $\alpha=0.2,\beta=5$, and (D) matching induced by TpOT with $\alpha=0.8,\beta=5$.
Intuitively, the parameter $\alpha$ controls the relative contributions of the geometric distortion (measured in terms of a Gromov-Wasserstein loss, with coefficient $\alpha$) and topological distortion (measured in terms of transport cost on persistence features, with coefficient $1-\alpha$). When $\beta=0$, point clouds and persistence diagrams are matched independently. For $\beta>0$, the hypergraph distortion $(\pi^{v},\pi^{e})\mapsto\left\langle L(\omega_{H},\omega_{H^{\prime}}),\pi^{v}\otimes\pi^{e}\right\rangle$ couples the matching of the point clouds and persistence diagrams, allowing geometric information to inform the topological matching and vice versa. In the limit as $\beta\to\infty$, this term dominates the overall objective in (16) and in the limit we recover a partial matching variant of HyperCOT [11] in which hyperedges can be transported to a null edge (corresponding to the diagonal in the persistence diagram setting) for a cost equal to the squared $L^{2}$-norm of the corresponding incidence function.
Theorem 2 allows us to explicitly construct a geodesic $[P_{t}],t\in[0,1]$ between $P_{X}$ and $P_{X^{\prime}}$.
@@
At each time $t\in[0,1]$, the corresponding gauged measure space $P_{t}$ can be represented as a point cloud with a gauge function $\omega_{t}$ given by interpolation of $\omega$ and $\omega^{\prime}$. In order to visualise the family of interpolating points, we consider the function $C_{t}((x,x^{\prime}),(y,y^{\prime}))\mapsto(1-t)C(x,y)+tC^{\prime}(x^{\prime},y^{\prime})$ where $C(x,y)$ and $C^{\prime}(x^{\prime},y^{\prime})$ are the squared Euclidean distances on the source and target point clouds respectively; this amounts to linear interpolation of $C$ and $C^{\prime}$ in $L^{2}_{\pi^{v}\otimes\pi^{v}}((X\times X^{\prime})^{2})$. At each value of $t$, positions of points in the interpolating point cloud are obtained by applying multidimensional scaling (MDS) algorithm to $C_{t}$ followed by an alignment step to remove issues due to the invariance of MDS under rigid transformation. Figure 3(E) shows a snapshot of this geodesic computed numerically for $\alpha=0.5,\beta=1$. We note that the geodesic almost perfectly recovers a homeomorphisms connecting each loop with their matched petal.
@@
Figure 5: (A) Source point cloud: a noisy mug in $\mathbb{R}^{3}$. (B) Target point cloud, a noisy solid torus in $\mathbb{R}^{3}$. Matching induced by GW distance. The colour code represents the matching. (C) Target point cloud, matching induced by TpOT with $\alpha=0.1$ and $\alpha=1$. The colour code represents the matching. (D) Persistence diagrams for source (blue) and target (yellow) point clouds. Both diagrams contain a unique significant homology class.
By construction of $X$, representative cycles for each of the four persistent homology classes span the circles almost entirely, and similarly for $X^{\prime}$. This makes the example in Figure 3 a relatively simple one, as most of the points in $X$ and $X^{\prime}$ are part of “topologically significant” cycles. We next investigate what happens for noisy point clouds which contain topological features with short lifetimes, or “topological noise”. Figure 4 shows an example of transport between point clouds containing noisy disks, and in this case the significant features in each of the persistent diagrams all have different lifespans. For small values of $\alpha$ (e.g. $\alpha=0.2$, Figure 4(C)) the influence of the topological distortion in the PD-space is greater. As a result, TpOT matches points in cycles with similar lifespan. On the other hand, for higher values (e.g. $\alpha=0.8$, Figure 4(D)) the geometric distortion has greater influence, and we observe that the matching tends to preserve local proximity between loops, rather than their relative sizes.
@@
Figure 6: (A) A point cloud $X$ consisting of 3 noisy loops of various sizes. (B) A point cloud $X$ consisting of the same 3 noisy loops as $X$, placed in different ways. Points forming the three persistent cycles in $\mathcal{K}_{\rm VR}(X)$ and $\mathcal{K}_{\rm VR}(X)$ are coloured in blue, green, and yellow. (C) Persistence diagrams of $\mathcal{K}_{\rm VR}(X)$ and $\mathcal{K}_{\rm VR}(X)$ plotted super-imposed to each other and showing Wasserstein (left) and Geometric (right) matchings. The three highly persistent classes are coloured as the corresponding cycles in (A) and (B). Geometric and Wasserstein matchings are shown with arrows in (A) and (B) as well. (D) A point cloud $X_{1}$ consisting of a chain of 9 noisy circles. (E) Persistence diagrams of $X_{1}$ super-imposed to that of $X_{2}$ (point cloud in (F-G)). Colour indicates points forming the 9 second-most persistent cycles in each point cloud. (F) The result of cycle matching induced by Wasserstein distance. (G) The result of geometric cycle matching. Colour indicates matched cycles.
@@
A more challenging example is shown in Figure 5. Here the objective is to transport a noisy point cloud $M$ resembling a mug (see Figure 5(A)), into a noisy solid torus $T$ (see Figure 5(B)) whilst preserving the topological features. In this context, matching of persistent homology features amounts to mapping the “handle” in $M$ into an essential (i.e. not bounding a disk) closed curve in $T$. As shown in Figure 5(D), the persistence diagrams of $T$ and $M$ have only one significant class each. That implies that outside of the points involved in the two chosen representative cycles, points in $M$ and $T$ have almost trivial topological signal. For this example, we set $\beta=5$. For low values of $\alpha$, the $\pi^{v}$ matches points successfully maps the handle in an essential closed curve spanning the doughnut hole (Figure 5(C), left), while for $\alpha=1$ (Figure 5(C), right) this effect is lost. This can be explained by looking at the matching between cycles induced by $\pi^{e}$, Figure 5(D): for low values of $\alpha$, the Wasserstein component promotes matching the highly persistent classes together, while for $\alpha=1$ these are split into less significant classes.
Figure 7: (A) Snapshots of a trefoil undergoing thermal relaxation. (B) Top rows shows the trefoil at step 4 of the simulation, while mid and bottom rows show the trefoil at step 5. Each column correspond to one of the most persistent homology classes of the trefoil at step 4. Mid and bottom rows shows the matched class at step 5, according to Wasserstein distance (mid) and TpOT (bottom). (C) Box plot of correlation values between cycles and their images.

### v2

@@ 4.3 Geodesics in the space of measure topological networks
$=\|\omega_{t}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{t}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{t},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}$
$\quad+t(1-t)\left(\|\omega_{0}-\omega_{1}\|^{2}_{L^{2}(\pi^{v}\otimes\pi^{e})}+\|k_{0}-k_{1}\|^{2}_{L^{2}(\pi^{v}\otimes\pi^{v})}+\|d_{2}\circ(\iota_{0},\iota_{1})\|^{2}_{L^{2}(\pi^{e})}\right)$
$=\|\omega_{t}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{t}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{t},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}$
$\quad+t(1-t)\left(\|\omega_{0}-\omega_{1}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{0}-k_{1}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{0},\iota_{1})\|^{2}_{L^{2}(\xi^{e})}\right)$ (41)
$=(1-t)\left(\|\omega_{0}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{0}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{0},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}\right)$
$\quad+t\left(\|\omega_{1}-\omega^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{e})}+\|k_{1}-k^{\prime}\|^{2}_{L^{2}(\xi^{v}\otimes\xi^{v})}+\|d_{2}\circ(\iota_{1},\iota^{\prime})\|^{2}_{L^{2}(\xi^{e})}\right)$ (42)
$\geq(1-t)d_{\rm TpOT,2}(P^{0},P^{\prime})^{2}+td_{\rm TpOT,2}(P^{1},P^{\prime})^{2},$ (43)
where (41) follows by marginal properties of $\xi^{v}$ and $\xi^{e}$, (42) by definition of $w_{t},k_{t}$ and $\nu_{t}$, and (43) by sub-optimality. ∎
5 Numerical examples
In this section, we demonstrate our computational framework and its capabilities on a range of numerical examples. In what follows, persistent homology computations are performed using the Julia package Ripserer.jl [16] with the Vietoris-Rips filtration. We emphasise that the choice of filtration, representatives, and construction of the hypergraph incidence function $\omega$ influences the resulting transport plan and analysis.
Figure 3: (a) Source point cloud $X$, (b) (i) Pairwise affinity matrix $C$ encoding geometric information; (ii) PH-hypergraph incidence matrix $\omega$ encoding the membership of points to persistence features; (iii) 1-dimensional persistence diagram $D$, points coloured by lifespan; (iv) columns of $\omega$ for the top 4 persistence features. Node-level features in $\omega$ are computed using TOPF [31], keeping the top 10 features. (c) Target point cloud $X^{\prime}$. (d) Same as (b). (e) TpOT ($\alpha=0.5,\beta=1$) and GW matching. Inset: matching $\pi^{e}$ of the persistence diagrams $D,D^{\prime}$ as found by TpOT. (f) A geodesic between the point clouds.
@@ 5.1 Matching point clouds
In Figure 3, we consider a source point cloud $X$ consisting of four noisy circles of uniform size (Figure 3(a)) and a target point cloud $X^{\prime}$ resembling a “flower” with four noisy ellipses as “petals” (Figure 3(c)). The point clouds $X$ and $X^{\prime}$, together with the uniform distribution on points, are then measure spaces. We choose to represent geometric information by taking the affinity matrices (also referred to as gauge functions) to be Gaussian kernels $C,C^{\prime}$, i.e. $C_{ij}=K(x_{i},x_{j})$ and $C^{\prime}_{ij}=K(x_{i}^{\prime},x_{j}^{\prime})$ where
$K(x,y)=\exp\left(-\frac{\|x-y\|_{2}^{2}}{h^{2}}\right).$
The bandwidth $h$ is chosen for each set of points $\{x_{1},\ldots,x_{N}\}$ such that $h^{2}N^{-2}\sum_{ij}\|x_{i}-x_{j}\|_{2}^{2}=1$.
@@
We compute $1$-dimensional persistent homology using the Vietoris-Rips filtration, and keep the top 10 features by persistence. We find that $\mathcal{K}_{\rm VR}(X)$ contains 4 significant, almost identical, homology classes (one per circle), and the same is true for $\mathcal{K}_{\rm VR}(X^{\prime})$ (one class per ellipse). A choice of generating cycle for each class in the diagrams yields the PH-hypergraphs $H_{X}=H(X,\mathcal{K}_{\rm VR}(X),g_{X})$ and $H_{X^{\prime}}=H(X^{\prime},\mathcal{K}_{\rm VR}(X^{\prime}),g_{X^{\prime}})$ and the corresponding topological networks (see Section 3.3)
$P_{X}$ $=\left((X,C,\mu),(Y,\iota,\nu),\omega\right)$
$P_{X^{\prime}}$ $=\left((X^{\prime},C^{\prime},\mu^{\prime}),(Y^{\prime},\iota^{\prime},\nu^{\prime}),\omega^{\prime}\right).$
As we point out earlier, the choice of pointwise representation of persistence features can affect results greatly. To improve robustness to noise, we employ the TOPF approach of [31] to compute the PH-hypergraph membership function for each persistence feature. In Figure 3(b) and (d), we see, from left to right: the target and source point clouds, their pairwise affinity matrix, the PH-hypergraphs incidence matrices $\omega,\omega^{\prime}$, their persistence diagrams, and a representation of their representative cycles as encoded in the corresponding PH-hypergraph. Using the algorithms we developed in Section 3.3, we are able to simultaneously find matchings between points and topological features in $P_{X}$ and $P_{X^{\prime}}$ as the parameters $(\alpha,\beta)$ vary.
We compute TpOT matchings of $P_{X}$ and $P_{X^{\prime}}$ using Algorithm 1 with $\varepsilon_{v}=3\times 10^{-3}$, $\varepsilon_{e}=10^{-2}$, $\alpha=0.5$, $\beta=1.0$. For comparison, we compute the (unregularised) Gromov-Wasserstein matching of $X$ an $X^{\prime}$ using the PythonOT package [24].
In Figure 3(e), points are coloured by their corresponding source points under the TpOT and Gromov-Wasserstein matchings respectively. From this we observe that topological features (i.e. closed loops) fail to be preserved by the GW matching. Instead, each of the circles in the source point cloud is split up among multiple ellipses in the target point cloud. The TpOT matching, on the other hand, simultaneously preserves local adjacency and topological features.
Figure 4: (a) Source point cloud $X$ and its persistence diagram $D$. (b) Target point cloud $X^{\prime}$ and its persistent diagram $D^{\prime}$. (c) TpOT ($\alpha=0.5,\beta=1$) and GW matching. (d) Matching $\pi^{e}$ of the persistence diagrams $D,D^{\prime}$ found by TpOT. (e) The variation of TpOT matchings as the parameters $\alpha$ and $\beta$ vary.
Given couplings $(\pi^{v},\pi^{e})$, Theorem 6 allows us to explicitly construct a geodesic $[P_{t}],t\in[0,1]$ between $P_{X}$ and $P_{X^{\prime}}$. At each time $t\in[0,1]$, the corresponding gauged measure space $P_{t}$ can be represented as a point cloud with a gauge function $\omega_{t}$ given by interpolation of $\omega$ and $\omega^{\prime}$. In order to visualise the family of interpolating points, we consider the function $C_{t}((x,x^{\prime}),(y,y^{\prime}))\mapsto(1-t)C(x,y)+tC^{\prime}(x^{\prime},y^{\prime})$ where $C(x,y)$ and $C^{\prime}(x^{\prime},y^{\prime})$ are the squared Euclidean distances on the source and target point clouds respectively; this amounts to linear interpolation of $C$ and $C^{\prime}$ in $L^{2}_{\pi^{v}\otimes\pi^{v}}((X\times X^{\prime})^{2})$. Since in practice we use an entropic regularisation, the resulting couplings $(\pi^{v},\pi^{e})$ are a priori strictly positive for all entries. We use the following heuristic sparsifying step, which was also used by [33]: given a coupling $\pi_{\varepsilon}\in\Pi(\mu,\mu^{\prime})$, we “round” it onto an extreme point of the coupling polytope by solving
$\max_{\pi\in\Pi(\mu,\mu^{\prime})}\langle\pi_{\varepsilon},\pi\rangle,$
which in itself amounts to solving a discrete optimal transport problem. This yields a sparse coupling approximating the solution found by entropic regularisation.
@@
At each value of $t$, positions of points in the interpolating point cloud are obtained by applying the multidimensional scaling (MDS) algorithm to $C_{t}$, followed by an alignment step to remove issues due to the invariance of MDS under rigid transformations. Figure 3(f) shows a snapshot of this geodesic computed numerically for $\alpha=0.5,\beta=1$. We note that the geodesic almost perfectly recovers a local homeomorphism connecting each loop with their matched petal.
@@
Figure 5: (a) Source point cloud: a noisy mug in $\mathbb{R}^{3}$. (b) Target point cloud, a noisy solid torus in $\mathbb{R}^{3}$, with matching induced by TpOT and GW distance, respectively. (c) The variation of TpOT matching as the parameters $\alpha$ and $\beta$ vary.
In Figure 4 we consider two point clouds $X,X^{\prime}$ which have nearly indistinguishable persistence diagrams $D,D^{\prime}$. As previously, we construct measure topological networks $P_{X},P_{X^{\prime}}$ with Gaussian affinity matrices and using TOPF with a Vietoris-Rips filtration [31] for topological features. We compute the TpOT matchings as before (Algorithm 1) with $\alpha=0.5,\beta=1,\varepsilon_{v}=3\times 10^{-3},\varepsilon_{e}=10^{-2}$. The choice of parameters $\alpha$ and $\beta$ for balancing the weights for the TpOT terms (see Equation 17) is a key factor influencing the behaviour of the matching. Of the three terms, the key one is the “cross-term” $\left\langle L(\omega,\omega^{\prime}),\pi^{v}\otimes\pi^{e}\right\rangle$ corresponding to the co-cost function, since this determines the strength of coupling between the geometric and generator matchings. When this term is given a zero weight (i.e. when $\beta=0$), the TpOT problem simplifies to that of independently finding geometric and generator matchings using a Gromov-Wasserstein and persistence diagram cost, respectively. Figure 4 shows the effect of varying $\beta$ on the example of Figure 4. As $\beta$ increases, so does the strength of the coupling between these two terms. In the limit as $\beta\to\infty$, this term dominates the overall objective in (17). In the limit we recover a partial matching variant of HyperCOT [13], in which hyperedges can be transported to a null edge (corresponding to the diagonal in the persistence diagram setting) for a cost equal to the squared $L^{2}$-norm of the corresponding incidence function. For positive $\beta$, the parameter $\alpha$ controls the relative contributions of the geometric distortion (measured in terms of a Gromov-Wasserstein loss, with coefficient $\alpha$) and topological distortion (measured in terms of transport cost on persistence features, with coefficient $1-\alpha$). For small values of $\alpha$ (e.g. $\alpha=0.1$), Figure 4 the influence of the topological distortion in the PD-space is greater. As a result, TpOT matches points in cycles with similar lifespan. On the other hand, for higher values (e.g. $\alpha=0.9$) the geometric distortion has greater influence, and we observe that the matching tends to preserve local proximity between loops, rather than their relative sizes.
@@
Figure 6: (a) A point cloud $X$ consisting of 3 noisy loops of various sizes. (b) On the left-hand side: a point cloud $X^{\prime}$ consisting of the same 3 noisy loops as $X$, placed in different ways. Points forming the three persistent cycles in $\mathcal{K}_{\rm VR}(X)$ and $\mathcal{K}_{\rm VR}(X^{\prime})$ are coloured according to TpOT-induced geometric matching. On the right-hand side: same as on the left, but points forming the three persistent cycles in $\mathcal{K}_{\rm VR}(X)$ and $\mathcal{K}_{\rm VR}(X^{\prime})$ are coloured according to Wasserstein matching on persistent diagrams. (c) A point cloud $X$ consisting of a chain of 9 noisy circles, matched to a copy of itself $X^{\prime}$. Colour indicates cycle matching induced by TpOT-induced geometric matching and Wasserstein matching on persistent diagrams, respectively.
@@
We show a more challenging example in Figure 5, in which the objective is to transport a noisy point cloud $X$ resembling a mug (see Figure 5(a)), into a noisy solid torus $X^{\prime}$ (see Figure 5(b)) whilst preserving the topological features found by persistent homology. In this context, matching of topological features amounts to mapping of the “handle” in $X$ into an essential (i.e. not bounding a disk) closed curve in $X^{\prime}$. The persistence diagrams of $X$ and $X^{\prime}$ have only one significant class each. This implies that, outside of the points involved in the two chosen representative cycles, points in $X$ and $X^{\prime}$ have almost trivial topological signal. To construct measure topological networks $P_{X},P_{X^{\prime}}$, we construct the PH-hypergraph incidence functions $\omega,\omega^{\prime}$ using the following smoothing procedure applied to a binary point-generator incidence function $\omega_{\mathrm{binary}}$:
$\omega_{\mathrm{smooth}}=\min_{x}\frac{1}{2}\|x-\omega_{\mathrm{binary}}\|_{F}^{2}+\frac{\lambda}{2}\operatorname{tr}(x^{\top}Lx)=(I+\lambda L)^{-1}\omega_{\mathrm{binary}},$
where we take $\lambda=1$ and $L$ is a symmetrically normalised Laplacian operator constructed, in our case, from a Gaussian kernel on each point cloud. The pairwise affinity matrices $C,C^{\prime}$ are also constructed using Gaussian kernels as done for the examples in Figures 3, 4.
We compute the TpOT matchings using Algorithm 1 with $\varepsilon_{v}=10^{-2},\varepsilon_{e}=10^{-2}$, and various choices of the parameters $\alpha,\beta$. For $\beta=2.5$ and low values of $\alpha$ (e.g., $\alpha=0.1$), the coupling $\pi^{v}$ matches points successfully maps the handle in an essential closed curve spanning the doughnut hole (Figure 5(b), left). When $\alpha$ increases (Figure 5(c)), this effect is lost since the Gromov-Wasserstein component dominates. On the other hand, fixing $\alpha=0.1$ and allowing $\beta$ to vary, we find that the correspondence between generators is noisy for smaller values $\beta=1$ but becomes stronger as $\beta$ increases. This can be understood as the increasing contribution of the coupling between points and topological features.
Figure 7: (a) Snapshots of a trefoil undergoing thermal relaxation. (b) Box plot of correlation values between cycles and their images. (c) Step by step tracking of cycles via TpOT-induced geometric matching (left) and Wasserstein matching on persistence diagrams (right).